OpenStudy (anonymous):

Find the real zeros of the trigonometric function on the interval 0 ≤ x < 2π f(x) = 4 cos^2(x) - 3

OpenStudy (campbell_st):

well you have \[4\cos^2(x) - 3 = 0\] or \[4\cos^2(x) = 3\] just keep solving...

OpenStudy (anonymous):

cos(x)= √3/2

OpenStudy (anonymous):

answers are π/6, 11π/6 ?

OpenStudy (campbell_st):

I think you will find there are 4 zeros.... remember taking the square root means plus and minus root 3

OpenStudy (anonymous):

okay, π/6, 5π/6, 7π/6, 11π/6

OpenStudy (campbell_st):

looks better... well done